### v1

# SceneSplat++: A Large Dataset and Comprehensive Benchmark for Language Gaussian Splatting

Abstract
3D Gaussian Splatting (3DGS) serves as a highly performant and efficient encoding of scene geometry, appearance, and semantics. Moreover, grounding language in 3D scenes has proven to be an effective strategy for 3D scene understanding. Current Language Gaussian Splatting line of work falls into three main groups: (i) per-scene optimization-based, (ii) per-scene optimization-free, and (iii) generalizable approach. However, most of them are evaluated only on rendered 2D views of a handful of scenes and viewpoints close to the training views, limiting their ability and insight into holistic 3D understanding. To address this gap, we propose the first large-scale benchmark that systematically assesses these three groups of methods directly in 3D space, evaluating on 1060 scenes across three indoor datasets and one outdoor dataset. SceneSplat-Bench results demonstrate a clear advantage of the generalizable paradigm, particularly in relaxing the scene-specific limitation, enabling fast feed-forward inference on novel scenes, and achieving superior segmentation performance. We further introduce SceneSplat-49K, a carefully curated 3DGS dataset comprising around 49K diverse indoor and outdoor scenes obtained from multiple sources, with which we demonstrate the generalizable approach could harness strong data priors. Our codes, benchmark, and datasets will be made public to accelerate research in generalizable 3DGS scene understanding.
[…]
$\ast$ indicates equal contribution. $\dagger$ indicates the corresponding author: Bin Ren <bin.ren@insait.ai>.
1 Introduction
The field of 3D computer vision is highly focused on capturing the geometry WITKIN198117 ; Kazhdan2006 ; Hirschmuller2008StereoDepth ; Newcombe2011KinectFusion ; schoenberger2016sfm ; Park2019DeepSDF ; shi4568335city ; ren2024bringing and the visual appearance 10.5555/539405 ; 791235 ; 5226635 ; mildenhall2020nerf ; kerbl3Dgaussians of a scene, as well as understanding its content 8099499 ; wu2024ptv3 ; peng2023openscene ; zhu2023pointclip ; sarkar2025crossover . Recently, 3D Gaussian Splatting (3DGS) kerbl3Dgaussians has emerged as the most desirable 3D representation due to its unique ability to jointly encode geometry, appearance, and scene understanding properties Ye2024GSgroupping ; Shi2024LEG ; jiang2024open in a compact form that can be effectively optimized from 2D posed images ye2025gsplat . Furthermore, vision-language reasoning represents the most promising direction for 3D scene understanding rozenberszki2022language ; ding2022pla ; yang2024regionplc ; ma2025cityloc ; jiang2024open ; wu2023mars , as it bridges the visual and geometric attributes of a scene with the language we use to define, describe, and reason about concepts. Therefore, this work focuses exclusively on vision-language scene understanding with 3DGS.
The most relevant methods for language Gaussian Splatting (LGS) can be categorized into three groups. The first two groups begin by extracting 2D features from all training images using vision-language foundation models (e.g., CLIP Radford2021CLIP ). The first group then performs gradient-based per-scene optimization Qin2024LangSplat ; Zuo2024FMGS ; zhou2024feature3DGS ; Wu2024OpenGaussians ; QU2024GOI , assigning feature vectors to each 3D Gaussian primitive and optimizing them such that their renderings align with the corresponding 2D feature maps. The second group also operates on a per-scene basis but adopts an optimization-free approach joseph2024GWfeatureBP ; marrie2024ludvig ; cheng2024occam . Instead of iterative refinement, these methods directly lift 2D features to 3D primitives via weighted feature aggregation schemes. Lastly, we consider the generalizable approach. To date, only a single method has been trained on 2K indoor scenes li2025SceneSplat7k , learning to directly predict a vision-language feature vector for each 3D Gaussian primitive. Although 2D vision-language features are required during training, inference relies solely on 3D Gaussians, eliminating the need to prepare additional 2D feature maps. Furthermore, all features are produced in a single forward pass.
While these methods represent important advances in integrating vision-language reasoning with 3DGS, their evaluation protocols still suffer from several critical limitations. First, most methods are evaluated on only a small number of selected scenes Qin2024LangSplat ; Zuo2024FMGS ; zhou2024feature3DGS ; Wu2024OpenGaussians ; QU2024GOI ; joseph2024GWfeatureBP ; cheng2024occam . As a result, such evaluations carry a high risk of producing scene-specific rather than generalizable conclusions. Moreover, the absence of a standardized benchmark further limits systematic and comparable evaluation. Second, most methods are evaluated close to the training views Qin2024LangSplat ; marrie2024ludvig ; Wu2024OpenGaussians . Consequently, there is no guarantee that the results generalize well to novel viewpoints. Finally, most evaluations are conducted in 2D rather than rather than directly in 3D space Qin2024LangSplat ; Zuo2024FMGS ; zhou2024feature3DGS ; Wu2024OpenGaussians ; QU2024GOI ; joseph2024GWfeatureBP ; marrie2024ludvig ; cheng2024occam . However, 3D scene understanding fundamentally concerns performance in 3D space, which cannot be fully inferred from 2D projections. Therefore, we introduce SceneSplat-Bench, an evaluation benchmark for 3DGS vision-language scene understanding that includes 1,060 scenes spanning 325 unique semantic classes. It evaluates performance in 3D for each segment of the scene, and we use it to assess representative methods from all three groups, as presented in Tab. 1.
Among the evaluated approaches, the generalizable 3D scene understanding paradigm demonstrates the strongest potential. It eliminates the need for substantial per-scene computation at inference and enables a single forward pass per 3D scene. In addition, this approach enables the application of 3D foundation models for scene understanding without the need for task- or scene-specific retraining. This concept aligns with established practices in 2D computer vision, where pretrained foundation models have been widely adopted He_2016_CVPR ; dosovitskiy2020vit ; oquab2023dinov2 ; ravi2024sam2 ; Radford2021CLIP ; liu2023llava . Foundation models tailored to 3DGS have already emerged across key areas, including reconstruction charatan23pixelsplat ; chen2024mvsplat ; gslrm2024 ; ziwen2024llrm , scene understanding li2025SceneSplat7k , and generation xiang2024structured ; roessle2024l3dg . Lastly, benchmark results demonstrate the state-of-the-art performance of the generalizable approach, highlighting its capacity to leverage data priors and extract meaningful vision-language features—effectively mitigating the noise inherent in weak supervision.
To facilitate the development of generalizable 3DGS scene understanding, we introduce SceneSplat-49K, a carefully curated 3DGS dataset comprising diverse indoor and outdoor scenes collected from multiple sources. Existing large-scale 3DGS datasets primarily focus on single objects ma2024shapesplat ; liu24uco3d ; wei20253daffordsplatefficientaffordancereasoning , whereas scene-level datasets li2025SceneSplat7k remain limited in scale. To the best of our knowledge, SceneSplat-49K represents the most extensive open-source dataset of complex, high-quality 3DGS reconstructions at the full-scene level. The preparation of the dataset required approximately 891 GPU days and considerable human involvement. Furthermore, we demonstrate that training a generalizable 3DGS scene understanding method on a larger subset from SceneSplat-49K leads to improved performance, achieving state-of-the-art results on the benchmark.
Our contributions can be summarized as follows:
  • We introduce SceneSplat-Bench, the first benchmark for systematically evaluating vision-language scene understanding methods in the 3DGS domain.
  • We release SceneSplat-49K, a high-quality, carefully curated 3DGS dataset comprising 49K indoor and outdoor scenes.
  • We scale up the training and evaluation of a generalizable VLM for 3DGS-based scene understanding across both indoor and outdoor environments, offering new insights.
Methods Generalizable Optimization- free Runtime / Scene Benchmark Dataset Scene Number Class Number
Per-Scene Optimization Methods
LangSplat Qin2024LangSplat 621 min LERF kerr2023lerf , 3DOVS liu2024weaklysupervised3dopenvocabulary 9 88
FMGS Zuo2024FMGS 76 min LERF kerr2023lerf , 3DOVS liu2024weaklysupervised3dopenvocabulary 11 123
Feature3DGS zhou2024feature3DGS 235 min Replica straub2019replica 4 7
OpenGaussian Wu2024OpenGaussians 297 min LERF kerr2023lerf , ScanNet dai2017scannet (10scenes) 14 76
GOI QU2024GOI 252min Mip-NeRF360 barron2021mipnerf , Replica straub2019replica 8 30
Per-Scene Optimization-Free Methods
Gradient-Weighted 3DGS  joseph2024GWfeatureBP 4 min LERF kerr2023lerf , 3DOVS liu2024weaklysupervised3dopenvocabulary 9 88
LUDVIG marrie2024ludvig 5.6 min LERF kerr2023lerf , SPIn-NeRF spinnerf , NVOS ren2022neural 21 80
OccamLGS cheng2024occam 3.2 min LERF kerr2023lerf , 3DOVS liu2024weaklysupervised3dopenvocabulary 9 88
Generalizable Method
SceneSplat li2025SceneSplat7k 0.24 min ScanNet dai2017scannet , ScanNet++ yeshwanth2023scannet++ ,Matterport3D chang2017matterport3d 732 321
SceneSplat-Bench ScanNet dai2017scannet , ScanNet++ yeshwanth2023scannet++ , Matterport3D chang2017matterport3d , HoliCity zhou2021holicity 1060 325
Table 1: Language Gaussian Splatting (LGS) Benchmark Overview. Left: Grouped methods and their properties. Right: Benchmark characteristics. Our proposed SceneSplat-Bench benchmark evaluates the LGS methods at scale, across three indoor datasets and one outdoor dataset.
2 Related Works
Gaussian Splatting Datasets. Most existing 3DGS datasets remain object-centric, synthetic, or limited in scale and scope. Moreover, recent works charatan23pixelsplat ; roessle2024l3dg ; ziwen2024llrm ; chen2024splatformer showcase 3DGS generalization, yet do not fully release their datasets, thus limiting reproducibility and scalability. Among publicly available datasets, ShapeSplat ma2024shapesplat transforms 65K object-level CAD models into 3DGS. 3DAffordSplat wei20253daffordsplatefficientaffordancereasoning focuses on functionality, providing 23K object instances labeled with affordances across 21 categories. uCO3D liu24uco3d expands the object-centric landscape with over 1,000 annotated categories and 360 coverage. Furthermore, scene-level 3DGS remains underexplored, with SceneSplat li2025SceneSplat7k being the only available dataset providing 6.8K indoor scenes to support open-vocabulary 3D understanding, though its scope is limited to the indoor domain. In summary, existing datasets fall short of enabling generalizable, multi-modal learning across diverse scenes. To address this, we introduce SceneSplat-49K, a large-scale dataset of 49K 3DGS scenes spanning indoor and outdoor environments, enriched with vision-language embeddings to support scene-level understanding.
Open-Vocabulary Scene Understanding. Recent advancements in open-vocabulary models have significantly expanded the capabilities of vision-language understanding. Foundation models zhang2022dino ; oquab2023dinov2 ; kirillov2023segment ; ravi2024sam ; radford2021learning ; zhai2023sigmoid ; tschannen2025siglip enabled visual feature extraction at scale, effectively supporting tasks like detection, segmentation and open-vocabulary understanding. Recent works have also introduced 2D foundation models into 3D by utilizing NeRF mildenhall2021nerf or 3DGS kerbl20233d . Existing 3DGS approaches fall into three categories: per-scene optimization-based zhou2024feature3DGS ; zheng2024gaussiangrasper ; Qin2024LangSplat ; guo2024semantic ; QU2024GOI ; Zuo2024FMGS ; peng2023openscene ; Wu2024OpenGaussians , per-scene optimization-free joseph2024GWfeatureBP ; marrie2024ludvig ; cheng2024occam and generalizable li2025SceneSplat7k . Per-scene optimization-based LGS emerged first, with LERF kerr2023lerf ; rashid2023language and LangSplat Qin2024LangSplat integrating 2D foundation model features into NeRF- and GS-based models, a design adopted by many later methods. Per-scene optimization-free LGS, such as OccamLGS cheng2024occam , LUDVIG marrie2024ludvig and Gradient Weighted 3DGS joseph2024GWfeatureBP further optimized this process by employing feature lifting, directly projecting all 2D features into 3DGS, reducing computation time from hours to seconds. SceneSplat li2025SceneSplat7k is the first generalizable LGS method, trained and evaluated directly on 3D data, achieving significant gains in open-vocabulary scene understanding with substantially faster inference on novel scenes. However, most of these methods have been evaluated in constrained 2D settings, with a limited number of scenes and semantic classes. We therefore introduce a comprehensive benchmark and use it to systematically evaluate recent language Gaussian Splatting methods.
3D Generalizability. Large-scale training and generalization in 2D have become increasingly mature Radford2021CLIP ; tschannen2025siglip ; lyu2025realrag ; lyu2024unibind ; ravi2024sam ; ravi2024sam2 ; zhong2025omnisam . In contrast, achieving semantic and language generalization in 3D remains challenging and relatively underexplored chen2024splatformer ; PointMAE ; zhao2021point ; wu2022point ; wu2024ptv3 ; lee2025mosaic3d ; wu2025sonata . One major obstacle is the limited quantity and fragmented nature of 3D datasets, as no single dataset provides the needed combination of (i) large-scale coverage, (ii) multi-scale diversity (from object-centric to indoor and outdoor scenes), and (iii) rich multimodal information (geometry, appearance, and semantics). As a result, even state-of-the-art 3D models struggle to generalize broadly. Mosaic3D-5.6M compiles over 30K RGB-D scenes with 5.6 million mask-text pairs by leveraging 2D vision-language models, yielding an significant resource for open-vocabulary 3D segmentation lee2025mosaic3d . The SceneSplat-7K dataset supports open-vocabulary training but with limited scale li2025SceneSplat7k . These efforts represent early steps toward 3D foundation models, aiming to enable robust, optimization-free, and generalizable 3D reasoning across vision-language tasks. Therefore, we introduce SceneSplat-49K, a large and comprehensive dataset designed to unlock scaling laws in generalizable 3DGS tasks.
3 Dataset
Property SceneSplat-7K DL3DV-10K HoliCity Aria-ASE Crowdsourced Total
Data Statistics
Raw Scenes 9114 10K 6 253 25K 603 48 856 $\sim$49K
GS Scenes 7 916 6 376 5 940 25K 603 45 835 $\sim$46K
Data Type Indoor Indoor/Outdoor Outdoor Indoor Indoor/Outdoor Indoor/Outdoor
RGB Frames 4.72M 3.09M 48K 14.79M N/A 22.65M
Avg. GS / Scene 1.42M 1.48M 800K 1.27M 1.02M 1.26M
3DGS GPU Time [h] 3.59K 2.1K 0.79K 6.25K N/A 12.73K
Scenes w/ VLM Embedding 6 121 - 5 940 - - 12 061
Embedding GPU Time [h] 8 161 - 495 - - 8 656
Storage [TB] 2.76 2.91 0.26 2.2 0.23 8.36
Total 3DGS 11.27B 9.42B 4.75B 3.18B 615M 29.24B
Evaluation Metrics (Avg.)
PSNR [dB] $\uparrow$ 29.64 25.70 29.85 27.32 N/A 27.83
Depth Loss [m] $\downarrow$ 0.035 - 0.010 0.082 N/A 0.061
SSIM $\uparrow$ 0.897 0.845 0.921 0.906 N/A 0.898
LPIPS $\downarrow$ 0.212 0.203 0.122 0.230 N/A 0.209
[…]
Figure 2: Appearance, Geometry, and Scale Statistics of the SceneSplat-49K Dataset. Distributions of photometric (PSNR, SSIM, LPIPS) and geometric (depth $\ell_{1}$) reconstruction errors show consistently high-quality renders across scenes, while the wide spread in total Gaussian number and indoor/outdoor scene floor area demonstrates the dataset’s diversity. The curves are convolved from the bucket values and vertical dotted lines mark the mean of each metric.
3.1 Data Collection and Processing
To ensure high-quality 3DGS scenes, we implement several quality control measures throughout the optimization pipeline. Starting with the training views, we select scenes with at least 400 frames to ensure sufficient multi-view coverage. When depth information is available, we initialize the Gaussians at fused point cloud locations and apply depth supervision to maximize geometry quality. We use gsplat ye2025gsplat for 3DGS optimization. We filter blurry frames using the variance of the Laplacian as a sharpness metric. To efficiently compress the 3DGS scene, we employ the Markov Chain Monte Carlo strategy kheradmand20253d and add opacity and scale regularization. Once optimized, the dataset supports filtering 3DGS scenes based on PSNR and depth quality before using them as inputs for model training. We introduce the processing of each data source in the following.
SceneSplat-7K contains 7,916 curated 3DGS indoor scenes derived from ScanNet dai2017scannet , ScanNet++ yeshwanth2023scannet++ , Replica straub2019replica , Hypersim roberts2021hypersim , 3RScan wald2019rio , ARKitScenes baruch2021arkitscenes , and Matterport3D chang2017matterport3d .We further enrich over 6,000 of these scenes with vision–language embeddings to support open-vocabulary scene queries.
[…]
DL3DV-10K dataset covers 65 everyday environments, featuring both indoor and outdoor locations. It includes 10,510 high-quality videos captured by mobile devices and drones, along with sparse COLMAP reconstructions. For 3DGS training, we use an evenly sampled set of every 10 views for novel view evaluation, while the remaining views are used for training.
[…]
Crowdsourced Data. To diversify our dataset, we collected high-quality 3DGS scenes from various sources, including Polycam and Sketchfab. We have also distributed our questionnaire to the community to collect data.
3.2 Vision-Language Embedding Collection
Vision-language embeddings are extracted from the frames used to optimize the 3DGS. Unlike existing pretraining methods that align 3D primitives with text embeddings from vision-language models, we align each Gaussian directly in the image-embedding space, thereby preserving richer latent semantics. This approach avoids the information loss inherent in textual descriptions.
We follow the fusion strategy in li2025SceneSplat7k and employ a dynamic weighting mechanism that adaptively combines three distinct features: global context from the entire frame, local features with background, and masked features without background. This mechanism automatically balances contributions based on each segment’s contextual relationship—emphasizing background-inclusive features for integrated objects (e.g., keyboard with monitor) and object-specific features for isolated objects. This adaptive approach provides more nuanced semantic understanding than fixed weighting strategies.
4 Benchmark
We formulate our benchmark as follows. A scene is represented by a set of $N$ Gaussian primitives $\mathcal{G}=\{G_{j}\}_{j=1}^{N}$, where each $G_{j}\in\mathbb{R}^{59}$ encodes its position, covariance, color, and opacity. For every primitive we compute a language embedding $E_{j}^{G}\in\mathbb{R}^{d}$, obtained by per-scene processing or by inference with a learned encoder. The user provides a set of free-form text queries $C=\{c_{i}\}_{i=1}^{n}$. A frozen text encoder $P_{T}$ (e.g., CLIP) maps each query to an embedding $E_{i}^{T}=P_{T}(c_{i})$. Each Gaussian is labeled with the query whose embedding is most similar to its own:
$\text{label}(G_{j})\;=\;\arg\max_{i}\;\operatorname{cos}\bigl{(}E_{j}^{G},\,E_{i}^{T}\bigr{)}.$ (1)
This label field assigns every primitive to the most relevant query.
4.1 Benchmark settings
Evaluation Protocols. We evaluate 9 methods from 3 categories on our SceneSplat-Bench benchmark. We report two key metrics in 3D: foreground mean Intersection over Union (f-mIoU) and foreground mean accuracy (f-mAcc), which address object size imbalances and reduce viewpoint dependency compared with 2D evaluation. We unified the evaluation pipeline by first converting the language fields into a unified format $\mathcal{E}_{G}=\{E_{j}\}_{j=1}^{N}$, then computing the similarity with the queried text embedding and selecting the pair with the highest similarity. K-nearest neighbors are used to vote predictions from Gaussian splats to ground truth label positions. We exclude floor, wall, and ceiling classes as background for indoor scenes and the sky class for outdoor scenes. All runtime measurements were obtained on NVIDIA RTX A6000 GPUs, except for FMGS, whose training requires 50–80 GB of GPU memory; its runtime was therefore measured on NVIDIA H200 GPUs.
Benchmark Dataset. We evaluate the methods’ performance on indoor scenes using ScanNet (with 20 and 200 classes), ScanNet++ (100 classes), and Matterport3D (21 classes). Queries cover a wide range of object granularity—from small items like a mouse to larger objects such as a bed—reflecting the method’s capability to varying scales understanding. HoliCity (4 classes) is employed to evaluate the performance on outdoor scenes, emphasizing large structures under complex surrounding environment. Performance results are reported on randomly sampled 10-scene subsets for each of the four benchmarks across all methods. For methods demonstrating strong efficiency, we conduct a full evaluation on the complete validation set.
Baseline Methods. We categorize the methods based on their optimization paradigm and generalizability to novel scenes. (i) Per-Scene Optimization-based Methods assign learnable attributes to Gaussian splats and optimize them via backpropagation. However, rendering additional features with alpha compositing is computationally expensive and requires per-scene optimization. (ii) Per-Scene Optimization-Free methods use a training-free approach, lifting 2D features into 3D space as a weighted sum of 2D representations, requiring only a single forward rendering pass. (iii) Generalizable Method SceneSplat trains a feed-forward 3DGS encoder, learning to predict representations aligned with image embeddings. We further extend its training based on our dataset. SceneSplat (Pseudo Label) denotes the VLM embeddings used for its vision-language pretraining. We refer to the supplement for details on running each baseline.
Figure 3: Qualitative Results of Zero-Shot 3D Semantic Segmentation. The semantic classes "bicycle" and "kitchen table" are highlighted, which are not labeled in Ground Truth.
Method ScanNet20 (10 scenes) ScanNet20 (312 scenes) ScanNet200 (10 scenes) ScanNet200 (312 scenes)
f–mIoU f–mAcc f–mIoU f–mAcc f–mIoU f–mAcc f–mIoU f–mAcc
Per-Scene Optimization Methods
LangSplat 0.0153 0.0741 - - 0.0057 0.0106 - -
OpenGaussian 0.1186 0.2340 - - 0.0545 0.1086 - -
Feature-3DGS 0.1711 0.2512 - - 0.0541 0.0876 - -
FMGS 0.0916 0.2028 0.1169 0.2684 0.0641 0.1374 0.0579 0.1443
GOI 0.2295 0.4638 0.2115 0.4286 0.1193 0.2297 0.1001 0.2240
Per-Scene Optimization-Free Methods
Gradient-Weighted 3DGS 0.4184 0.5174 0.3573 0.4638 0.1638 0.2502 0.0601 0.1100
LUDVIG 0.1406 0.2885 0.1337 0.2969 0.0887 0.1833 0.0602 0.1701
OccamLGS 0.3559 0.4587 0.2308 0.4148 0.1936 0.2458 0.1204 0.2503
SceneSplat (Pseudo Label) 0.3521 0.5210 0.3500 0.5550 0.2520 0.3038 0.2280 0.3590
Generalizable Method
SceneSplat 0.2968 0.5711 0.3540 0.5780 0.1387 0.4198 0.1648 0.3573
Table 3: Zero-shot 3D Semantic Segmentation Experiments on ScanNet20 (20 classes) and ScanNet200 (200 classes). All methods are evaluated on a 10-scene mini-validation set, with the 312-scene full validation set run only for selected methods due to runtime limitations (Tab. 1).
Method ScanNet++ (10 scenes) ScanNet++ (50 scenes) Matterport3D (10 scenes) Matterport3D (370 scenes)
f–mIoU f–mAcc f–mIoU f–mAcc f–mIoU f–mAcc f–mIoU f–mAcc
Per-Scene Optimization Methods
LangSplat 0.0131 0.0332 - - 0.0219 0.1257 - -
OpenGaussian 0.0695 0.1186 - - 0.1387 0.2698 - -
Feature-3DGS 0.0798 0.1765 - - 0.3043 0.4732 - -
FMGS 0.0984 0.1892 0.1004 0.2300 0.2105 0.3664 - -
GOI 0.1593 0.2790 0.1631 0.3296 0.1631 0.3296 0.1652 0.3130
Per-Scene Optimization-Free Methods
Gradient-Weighted 3DGS 0.0925 0.1391 0.0900 0.1328 0.3083 0.4586 0.2762 0.4187
LUDVIG 0.0995 0.2099 0.1099 0.2547 0.2348 0.3888 0.1960 0.3738
OccamLGS 0.1580 0.2861 0.1502 0.3312 0.1803 0.3263 0.1725 0.3151
SceneSplat (Pseudo Label) 0.2340 0.3709 0.2243 0.4667 0.2392 0.4368 0.2748 0.4384
Generalizable Method
SceneSplat 0.2263 0.4916 0.2836 0.4992 0.2732 0.5379 0.3384 0.5745
Table 4: Zero-Shot 3D Semantic Segmentation Experiments On ScanNet++ and Matterport3D. All methods are evaluated on a 10-scene mini-validation set, with the full validation set evaluated only for selected methods due to runtime limitations (Tab. 1).
Training Source #Scene ScanNet++ ScanNet200 Matterport3D HoliCity
f–mIoU f–mAcc f–mIoU f–mAcc f–mIoU f–mAcc mIoU mAcc
Indoor Training-Data Scaling
ScanNet++ (v1) 280 0.1683 0.3230 0.1081 0.1650 0.1010 0.2108 0.1888 0.3003
ScanNet++ (v2) 906 0.2383 0.4324 0.1180 0.1920 0.1491 0.2756 0.2339 0.3524
ScanNet++ (v2) + ScanNet 2107 0.2779 0.5055 0.1427 0.3434 0.2289 0.3803 0.2633 0.3504
ScanNet++ (v2) + ScanNet + M3D 3503 0.2836 0.4992 0.1648 0.3573 0.3384 0.5745 0.2026 0.3600
Matterport3D only 1396 0.3244 0.5354 0.2240 0.3889
Outdoor-Domain Training
HoliCity only 3000 0.2880 0.6045
Table 5: Impact of Training-Data Scaling on Indoor Benchmarks and Cross-Domain Generalization to HoliCity. Results show that more training data consistently improves indoor performance. Furthermore, models trained on indoor data only surprisingly transfer to outdoor scenes.
     Method      10 scenes      val set, 328 scenes
     f–mIoU      f–mAcc      f–mIoU      f–mAcc
     Per-Scene Optimization Methods
     LangSplat      0.0918      0.1950      -      -
     OpenGaussian      0.1853      0.2567      -      -
     Feature-3DGS      0.2479      0.5646      -      -
     FMGS      0.1774      0.2513      -      -
     GOI      0.1126      0.3773      0.1587      0.3772
     Per-Scene Optimization-Free Methods
     Gradient-Weighted 3DGS      0.2484      0.4808      0.2762      0.4187
     LUDVIG      0.1593      0.2296      0.1843      0.2775
     OccamLGS      0.2255      0.4754      0.2234      0.5056
     SceneSplat (Pseudo Label)      0.2464      0.4780      0.2522      0.5159
     Generalizable Method
     SceneSplat      0.3081      0.5647      0.2880      0.6045
Table 6: Zero-Shot 3D Outdoor Segmentation On HoliCity Dataset. All methods are evaluated on a 10-scene mini-validation set, with the full validation set evaluated for selected methods.
4.2 Key Insights
The primary experimental results from our SceneSplat-Bench benchmark across all selected methods are presented in Tabs. 3 and 4 for indoor scenes and in Tab. 6 for outdoor scenes. The corresponding runtime statistics for individual methods can be found in Tab. 1. SceneSplat – representing the generalizable paradigm – stands out as the clear winner in terms of both performance and efficiency (see Fig. 3 for qualitative examples). Interestingly, in 75% of the experiments, it even outperforms its own pseudo labels used for pretraining (SceneSplat (Pseudo Label)). This result highlights its ability to generalize by leveraging large-scale data priors and learning to predict meaningful vision-language features, effectively mitigating the noise inherent in weak supervision. As illustrated in Fig. 4, visualizations of text query results using enhanced NeRFView tools nerfstudio ; yang2024spectrally ; li2025SceneSplat7k further demonstrate that the generalizable method outperforms the others. Moreover, unlike other methods that rely on extensive feature extraction, the generalizable approach requires only 3DGS as input at inference time and predicts per-primitive vision-language features in a single forward pass.
Among the per-scene methods, optimization-free approaches clearly outperform optimization-based ones, both in terms of segmentation accuracy, as shown in Tabs. 3, 4 and 6, and in runtime efficiency, as reported in Tab. 1. This is likely due to the fact that the objective function in optimization-based methods is typically designed to perform as well as possible on the training views, which may not be ideal to novel viewpoints. Taking this into account, along with the strong performance of the generalizable approach, we conclude that per-scene optimization is not necessary for effective vision-language reasoning in 3DGS. However, within both the optimization-based and optimization-free groups, no single method consistently dominates across all datasets, indicating that performance remains sensitive to dataset-specific characteristics.
A key requirement for reliable benchmarking is having a sufficient number of scenes and appropriately challenging evaluation settings. As shown in Tabs. 3, 4 and 6, the performance of various methods varies noticeably when scaling from a small subset of 10 scenes to the full validation set, highlighting the importance of the benchmark scale in assessing performance. Furthermore,  Tab. 3 demonstrates a clear drop in accuracy when the task complexity increases from 20 to 200 semantic classes, emphasizing the need for demanding benchmarks to reveal the limitations of competing approaches.
Tab. 5 presents our scaling study, which offers two key takeaways. First, training-data scaling consistently improves performance on indoor benchmarks: expanding the training set from 280 to 3503 scenes lifts ScanNet++ f-mIoU from 0.168 $\rightarrow$ 0.284 and ScanNet200 from 0.108 $\rightarrow$ 0.165. A similar effect appears when the model jointly trained on three indoor datasets outperforms that trained on Matterport3D alone. Second—and more surprisingly—an indoor-only model transfers to outdoor scenes, but domain-specific data is still required to close the gap. Without ever seeing outdoor data, the largest indoor model reaches 0.263 mIoU on HoliCity, an improvement of 7.4% over the 280-scene baseline. This suggests that large-scale indoor training induces representations that partially generalize across the indoor–outdoor boundary, see Fig. 5 for zero-shot examples. Nevertheless, a model trained only on HoliCity still leads (0.288 mIoU), indicating that cross-domain capability ultimately needs outdoor supervision. These findings underscore the value of our large-scale indoor–outdoor 3DGS dataset as a powerful training resource.
Figure 4: Text-Based Scene Query. Given the prompt "These are fruits" to different LGS methods, the queried parts are highlighted in red.
Figure 5: Zero-Shot Predictions of Indoor-Trained SceneSplat on Outdoor Scenes. The results highlight the cross-domain capability. Color palette denotes buildings, roads, terrains, and trees.
5 Conclusion and Limitations
This work introduces the first comprehensive benchmark for 3DGS-based scene understanding methods, along with a large-scale 3DGS dataset of diverse indoor and outdoor scenes. Our evaluation demonstrates that generalizable approach consistently outperforms per-scene methods, establishing a new paradigm for scalable scene understanding through pretrained models. Despite these contributions, limitations remain. Our dataset scale, while substantial, could be further expanded. The outdoor benchmark is constrained by limited semantic classes, and not all 3DGS scenes include precomputed vision-language embeddings. We leave these as important directions for future work and hope our benchmark and dataset will be helpful to advance language-grounded 3DGS scene understanding.